OpenStudy (anonymous):

Can someone please explain how to solve this problem? Thanks. Determine whether y varies directly with x. If so, find the constant of variation k and write the equation. X Y 4 6.4 7 11.2 10 16 13 20.8

OpenStudy (anonymous):

no y does not varie as x

OpenStudy (anonymous):

x/y should = to x in all examples

OpenStudy (anonymous):

how does it not vary with x, how do i check it myself

OpenStudy (phi):

y varies directly with x means y= k x where k is some constant number. This is one of the simplest equations you will meet. if you have an (x,y) pair, you can replace the x and y in the equation and solve for k: for the first entry in your table, you have x=4 and y is 6.4 replace y with 6.4, and x with 4 in y = k x to get 6.4= k * 4 divide both sides by 4: 6.4/4 = k * 4/4 or k= 6.4/4 = 1.6 so your direct variation equation is y= 1.6 x if the other pairs in your table "work" then you can say they all have the same equation, and y varies directly with x (with k= 1.6) can you test the other (x,y) values in your table ?

OpenStudy (anonymous):

yes, just give me a moment please

OpenStudy (anonymous):

all x*1.6 equals the y values in the chart so y DOES directly vary with x, right?

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank you so much, i think its easy now :D

OpenStudy (phi):

yw

OpenStudy (anonymous):

sorry it was my calcuation error....i did a calcualtion error for 13 and 20.8

OpenStudy (anonymous):

i am very sorry